# A Smart Mnemonic Sounds like “Glue Tonic”: Mixing LLMs with Student Feedback to Make Mnemonic Learning Stick

### v1

@@ 7 Related Work
Mnemonic Generation: Mnemonic devices help users learn information, such as passwords Yang et al. (2016); Song et al. (2019), vocabulary Dundes (1961); Levin et al. (1992), and medical facts Ajayi et al. (2019); Leeds et al. (2020). There are several types of mnemonics, including song Hayes (2009); Werner (2018), acronym McCabe et al. (2013); Li et al. (2021), and keyword mnemonics Campos et al. (2004, 2011). We study keyword mnemonics, which link complex terms to simpler keywords.
The effort of manually writing mnemonics has led to mnemonic generation research. Early works use phonetic similarity and multi-score ranking to find keywords Savva et al. (2014); Anonthanasap and Leelanupab (2015); Anonthanasap et al. (2015), but these methods do not explain how the keyword is linked to the fact. Recent works prompt LLMs to produce these explanations Lee and Lan (2023); Wong and Wolf (2024). Conversely, we collect fine-tuning and preference datasets to train a mnemonic generator guided by real-world student feedback.
Human Preferences: Recent works align LLMs with preference data, which capture what humans prefer Stiennon et al. (2020). Alignment methods include reinforcement learning with reward models Christiano et al. (2017); Ziegler et al. (2019), selecting high-quality data Sanh et al. (2022); Zhou et al. (2024), and augmenting LM loss with reward models Yuan et al. (2023); Rafailov et al. (2024). Preferences have been used in sentiment generation Maas et al. (2011), summarization Völske et al. (2017), and dialogue safety Bai et al. (2022), but we are the first to study preferences in mnemonics.
Our work also follows recent efforts to measure issues in preferences, such as preference agreement Ethayarajh and Jurafsky (2022); Bansal et al. (2024) and annotator biases Peng et al. (2022); Wan et al. (2023). In contrast, we discern expressed and observed preferences and show expressed preferences do not capture what truly helps users. We are also similar to works using diverse human preferences Xue et al. (2023); Kirk et al. (2024), which we explore via Bayesian modeling Yang et al. (2024).
@@ 8 Conclusion
We design Smart, the first keyword mnemonic generator guided by student feedback. Smart is trained on new fine-tuning and preference datasets which will both be released. While curating data, we find low agreement in expressed preferences and our introduced observed preferences, showing that students cannot predict their learning. Combining expressed and observed preferences via DPO and Bayesian modeling yields a smaller, more efficient mnemonic model matching GPT-4. However, our human writer surpasses both models, especially in keyword simplicity and explanation quality, motivating mnemonic generation as a difficult task.
@@ 9 Limitations
One limitation is that our fine-tuning and preference datasets are relatively small. Despite this, our datasets both improve the quality of mnemonics from Smart (section 6.1), following the recent paradigm of LIMA Zhou et al. (2024) which suggests that small, high-quality datasets can be used to align and improve LLMs. Further, regardless of size, our preference dataset allows us to conduct an insightful analysis on the relation between expressed and observed preferences (section 4). As more students study in our app, we will update and release our preference datasets accordingly, resulting in more data to facilitate mnemonic research.
We also note that using GPT-4 as a judge can result in biases. Our decision to use GPT-4 was motivated by the high cost of human annotations, which has also become a standard practice Chiang et al. (2023); Liu et al. (2023); Touvron et al. (2023); Chiang and Lee (2023); Dettmers et al. (2023). While using GPT-4, we adopt several best practices for robust evaluation: 1) only evaluating on Smart to avoid self-recognition Panickssery et al. (2024); 2) using DSPy to help limit the sensitivity of prompt perturbations Khattab et al. (2024); 3) ensuring GPT-4 has high agreement with humans on held-out mnemonic comparisons Zheng et al. (2024); and 4) running inference on both orders of the mnemonic pairs to curb positional biases Wang et al. (2023b).
Lastly, our final model for Smart optimizes on a combination of multiple preference labels: pairwise comparisons, Likert ratings, and learning. While training and evaluating DPO models on all three objectives independently would be insightful, we are limited by our data. Thus, we focus on the most popular of the three—pairwise preferences—and our evaluation reveals how optimizing with multiple preference labels can impact metrics based on just one preference label. We encourage future works to use our preference data to explore how optimizing on multiple preferences impacts metrics based on each type of preference. To further motivate works in this direction, we have mnemonic experts evaluate Smart’s mnemonics to inform which types of feedback (e.g. keyword simplicity, explanation imageability) could be collected and trained with to improve mnemonics (section 6.3).
@@ 10 Ethical Considerations
While optimizing directly on observed preferences for downstream applications is promising, we advise researchers to take caution for deploying such models in a way that impacts users. We found that even our relatively harmless objective of optimizing on learning can result in bizarre mnemonics that perpetuate harms (Appendix D.1), and we imagine that in other domains, these consequences could be more severe. For example, training an LLM in the news domain optimized on user clicks can result in misinformation Milano et al. (2020), while optimizing directly on time spent looking at an advertisement can yield harmful, addictive content.
We took precautions to avoid these harms, such as filtering out offensive mnemonics before fine-tuning (section 2.3), and our protocols were approved by an Institutional Review Board to mitigate any harms in our user study. Thus, while we advocate for research in observed preferences for LLM alignment, we also urge researchers to consider the consequences of optimizing on these objectives before deployment in user-facing applications.
11 Acknowledgements
We would like to thank members of the CLIP lab at the University of Maryland and external collaborators for their discussions of this work, including David Martinez, Naina Balepur, Sandra Sandoval, Yoo Yeon Sung, and Rachel Rudinger. We also thank GregMat, GMAT Club, and Memli App for their feedback on our user study design and allowing us to use their educational materials in our research. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 2236417. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ A.2 $\mathcal{D}_{ft}$ and $\mathcal{D}_{pref}$ Dataset Details
In Tables 4 and 6, we provide descriptions of the columns in $\mathcal{D}_{ft}$ and $\mathcal{D}_{pref}$, respectively. For $\mathcal{D}_{pref}$, we filter out terms and mnemonics with less than two annotations, resulting in the summary statistics described in section 3.3, and summarized in Table 5. Along with the filtered subset, we will release the entire dataset of human preferences. We will also continue to release data as users study vocabulary with mnemonics in our app. The list of 500 vocab terms $\mathcal{V}_{test}$ used for testing will also be released.
Finally, in Table 7, we quantify the noise of annotator ratings in $\mathcal{D}_{pref}$. We use average entropy and variance as each instance in our dataset can be annotated by a different number of annotators, and there is no guarantee that the same annotator will be rating each of the mnemonics. We find that the average variance and entropy of our annotations is significantly lower than random chance.
Our datasets are based on publicly available GRE vocabulary words and mnemonics from MnemonicDictionary, and both were used as intended by the original authors. None of our datasets contain personal information, as users are referred to just by ID. All collected mnemonics are in English.
@@ B.4 Bayesian Model Evaluation
In this section, we evaluate the quality of our Bayesian model. Since our goal was to estimate the latent effectiveness of mnemonics, we first assess the convergence of our learned parameters across chains. All of our learned parameters have an $\hat{r}$ value under 1.01 and an effective sample size over 1000, indicating strong convergence. Further, our final Bayesian preference label (which mnemonic has a higher effectiveness score) across chains has a Krippendorff’s $\alpha$ over 0.75, indicating strong agreement and convergence. Finally, in Figure 9, we display the log likelihood values for our observed data across iterations, finding that they converge.
While convergence is more important to assess the quality of a Bayesian model that learns latent values, we also assess the generalizability of our model. We first train our model on 80% of our data and run inference on the remaining 20% for evaluation. In Figure 10, we compare the log likelihood of predicting the observed data on the training and evaluation splits. For 3/5 our observed data types, we find a non-significant difference between the log likelihoods (2-sample $t$-test). The only significant difference is in the data associated with the learning preferences $y_{learn}$, further suggesting that modeling observed preferences is a challenging and interesting direction for future work.
@@ B.5 Obtaining Bayesian Preference Labels
To obtain the final latent overall mnemonic effectiveness values for training the DPO model $p_{dpo}(m|v)$, we average the latent variables over all five chains post-burn-in. For the ablation study where we compare optimizing on $y_{pair}$ versus $y_{bayes}$, we similarly obtain the final latent mnemonic effectiveness values by averaging the latent variables over all five chains post-burn-in, but this time we just use a random sample of the epochs (i.e. thinning). By taking a sample of each chain instead of using the entire chains, we introduce more variability and disagreement in the labels, allowing us to more meaningfully study the difference between optimizing on pairwise preferences versus all preference labels.
@@ C.1 Annotator Instructions
During our user studies, we ensure to provide ample instructions to annotators. On the home page of our flashcard learning app, users can view our Institutional Review Board documents, which detail the purpose of the study and how user data will be collected and used. Further, we provide users instructions to help them rate the quality of mnemonic devices (Figure 7), which can be viewed at any time throughout the user study. All annotators are English speakers. Users were aware that they were participants in a research study and as participation was voluntary, compensation is fair.
@@ C.2 Qualitative Evaluation Details
In our qualitative evaluation, we compare our full model trained with DPO on Bayesian labels versus 10-shot GPT-4 (gpt-4-turbo-2024-04-09), where examples were chosen according to the highest latent quality scores in $\mathcal{D}_{train}$ (section 2.2). We also compare against expert human-written mnemonics to serve as an upper bound on mnemonic quality. These mnemonics were written by a professional copyeditor and creative writer with a Bachelor’s of Science degree that we hired on Upwork. As part of the interview, we asked the writer to produce two sample mnemonics to ensure the mnemonics would be high quality, and the annotator was paid a high rate of \$3 per mnemonic (around \$60/hr), which is fair for the participant demographic.
@@ D.2 Are Bayesian Tie-Breaks Good?
We previously found that using multiple preference labels to break ties within singular preference labels, improving LLM output quality with DPO (section 6.2). To ensure that these ties are better than random tie breaks, we use our GPT-4 classifier to compare the winning and losing mnemonics from our Bayesian labels (i.e. mnemonics with higher and lower effectiveness) where the majority vote in the pairwise setting is Tie. On the set of ties, GPT-4 states that our “winning” mnemonics is better 40% of the time, tied with the losing mnemonic 31% of the time, and is worse 29% of the time. Thus, even though humans marked these LLM outputs as tied, we were able to draw from other preference labels to identify a winning mnemonic, and GPT-4 also tends to agree that this mnemonic is higher-quality.
@@ D.3 DPO Models versus Fine-Tuning
In Table 11, we use GPT-4 to judge the mnemonics from each of the DPO configurations used in section 6.2 versus the Smart model just using fine-tuning $p_{0}(m|v)$. We find that DPO improves the mnemonic quality of each of these models.
@@ D.4 Mnemonic Examples
In Table 12, we provide examples of high-quality keyword mnemonic devices generated by our final model. We also show some examples of low-quality mnemonics from our model and highlight areas for improvement in Table 13.
@@
Figure 7: Instructions given to students to help them rate the quality of mnemonic devices for verbalized preferences.

### v2

@@ 7 Related Work
Below, we review relevant literature on mnemonic generation (section 7.1) and human preferences (section 7.2).
7.1 Mnemonic Generation
Mnemonics help users learn information, such as passwords Yang et al. (2016); Song et al. (2019), vocabulary Dundes (1961); Levin et al. (1992), and medical facts Ajayi et al. (2019); Leeds et al. (2020). There are many mnemonic types, including song Hayes (2009); Werner (2018), acronym McCabe et al. (2013); Li et al. (2021), and keywords Campos et al. (2004, 2011). We study keyword mnemonics, which link complex terms to simpler keywords.
The effort of manually writing mnemonics has led to mnemonic generation research. Early works use phonetic similarity and multi-score ranking to find keywords Savva et al. (2014); Anonthanasap and Leelanupab (2015); Anonthanasap et al. (2015), but these methods do not explain how the keyword is linked to the fact. Recent works prompt LLMs to generate mnemonic explanations Lee and Lan (2023); Wong and Wolf (2024), but we are the first to collect fine-tuning and preference data to generate mnemonics guided by real-world student feedback.
Prior education work has found that students cannot predict which study strategies, such as blocked versus interleaved practicing, will best help them learn McCabe (2011); Yan et al. (2016). We discover that students struggle to do the same for individual study items (section 4.2), such as mnemonics.
7.2 Human Preferences
Recent work aligns LLMs with preference data capturing what humans prefer Stiennon et al. (2020). Alignment methods include reinforcement learning with reward models Christiano et al. (2017); Ziegler et al. (2019), selecting high-quality data Sanh et al. (2022); Zhou et al. (2024), and augmenting LM loss with rewards Yuan et al. (2023); Rafailov et al. (2024). Preferences have been used for sentiment generation Maas et al. (2011), summaries Völske et al. (2017), and dialogue safety Bai et al. (2022), but we are the first to study them in mnemonics.
Our work also follows recent efforts to measure issues in preferences, such as preference agreement Ethayarajh and Jurafsky (2022); Bansal et al. (2024) and annotator biases Peng et al. (2022); Wan et al. (2023). In contrast, we distinguish between expressed and observed preferences and show expressed preferences do not capture what truly helps users. We are also similar to works that collect preferences across varied demographics Kirk et al. (2024) and use reinforcement learning regularization to align models to diverse preferences Xue et al. (2023), but we are the first to collect diverse preferences in an education setting and combine them via Bayesian modeling Yang et al. (2024).
In social science, several works find that a human’s stated preferences, elicited from survey responses, do not always agree with revealed preferences, the human’s actual behavior Urama and Hodge (2006); Hoderlein and Stoye (2014); De Corte et al. (2021). We show that the same applies in education, as pairwise comparisons and Likert ratings for perceived learning efficacy (expressed preferences) have low agreement (section 4.1) with what truly helps users learn (observed preferences). Mozannar et al. (2024) find a similar trend in coding, where programmer preferences do not correlate with their actual productivity. As a result, we hope our work will lead researchers to reevaluate how we should measure helpfulness in preference data collection to design models that truly help users downstream.
@@ 8 Conclusion
We design Smart, the first keyword mnemonic generator guided by student feedback. Smart is trained on new fine-tuning and preference datasets, both of which are released. While curating data, we reveal low agreement in expressed preferences and our introduced observed preferences, showing that students cannot predict their learning. Combining expressed and observed preferences via DPO and Bayesian modeling yields a smaller, more efficient mnemonic model matching GPT-4. However, our human writer surpasses both models, especially in keyword simplicity and explanation quality, motivating mnemonic generation as a challenging task. To further challenge LLMs, researchers could explore personalizing mnemonics for students, adapting Smart to different domains, languages, and modalities, or optimizing on other learning signals (e.g. long-term learning) via student feedback.
@@ 9 Limitations
One limitation is that our fine-tuning and preference datasets are relatively small. Despite this, our datasets both improve the quality of mnemonics from Smart (section 6.1), following the recent paradigm of LIMA Zhou et al. (2024) which suggests that small, high-quality datasets can be used to align and improve LLMs. Further, regardless of size, our preference dataset results in an insightful analysis of the relation between expressed and observed preferences (section 4). If more students study vocab in our app, we will update and release both of our mnemonic datasets accordingly, resulting in larger datasets to facilitate mnemonic research.
GPT-4 as a judge can result in biases. We used GPT-4 due to the high cost of human annotations, which has also become a standard practice Chiang et al. (2023); Liu et al. (2023); Touvron et al. (2023); Chiang and Lee (2023); Dettmers et al. (2023). While using GPT-4, we adopt best practices for robust evaluation: 1) only evaluating on Smart to avoid self-recognition Panickssery et al. (2024); 2) using DSPy to limit the sensitivity of prompt perturbations Khattab et al. (2024); 3) ensuring GPT-4 has high agreement with humans on held-out mnemonic comparisons Zheng et al. (2024); and 4) running inference on both orders of mnemonic pairs to curb position bias Wang et al. (2024).
Lastly, our final model for Smart optimizes on a combination of multiple preference labels: pairwise comparisons, Likert ratings, and learning. While training and evaluating DPO models on all three objectives independently would be insightful, we are limited by our data. Thus, we focus on the most popular of the three—pairwise preferences—and our evaluation reveals how optimizing with multiple preference labels can impact metrics based on just one preference label. We encourage future work to use our preference data to explore how optimizing on multiple preferences impacts metrics based on each type of preference. To further motivate works in this direction, we have mnemonic experts evaluate Smart’s mnemonics to inform which types of feedback (e.g. keyword simplicity, explanation imageability) could be collected and trained with to improve mnemonics (section 6.3).
@@ 10 Ethical Considerations
While optimizing directly on observed preferences for downstream applications is promising, we advise researchers to take caution for deploying such models in a way that impacts users. Even our relatively harmless objective of optimizing on learning can result in bizarre mnemonics that perpetuate harms (Appendix D.1), and we imagine that in other domains, these consequences could be more severe. For example, training an LLM in the news domain optimized on user clicks can result in misinformation Milano et al. (2020), while optimizing directly on time spent looking at an advertisement can yield harmful, addictive content.
We took precautions to avoid these harms, such as filtering out offensive mnemonics before fine-tuning (section 2.3), and our protocols were approved by an Institutional Review Board to mitigate any risks during our user study. Thus, while we advocate for research in observed preferences for LLM alignment, we also urge researchers to consider the consequences of optimizing on these objectives before deployment in user-facing applications.
Acknowledgements
We would like to thank members of the CLIP lab at the University of Maryland and external collaborators for their discussions of this work, including David Martinez, Naina Balepur, Sandra Sandoval, Yoo Yeon Sung, and Rachel Rudinger. We also thank GregMat, GMAT Club, and Memli App for their feedback on our user study design and for allowing us to use their testing materials in our research. This material is based upon work supported by the National Science Foundation under Grant No. iis-2403436 (Boyd-Graber) and dge-2236417 (Balepur). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Annotations and cloud computing resources were made possible by a gift from Adobe Research.
@@ A.2 $\mathcal{D}_{ft}$ and $\mathcal{D}_{pref}$ Dataset Details
In Tables 4 and 6, we provide descriptions of the columns in $\mathcal{D}_{ft}$ and $\mathcal{D}_{pref}$, respectively. For $\mathcal{D}_{pref}$, we filter out terms and mnemonics with less than two annotations, resulting in the summary statistics described in section 3.3, and summarized in Table 5. Along with the filtered subset, we will release the entire dataset of human preferences. We will also continue to release data if users study vocabulary with mnemonics in our app. The list of 500 vocab terms $\mathcal{V}_{test}$ used for testing will also be released.
Finally, in Table 7, we quantify the noise of annotator ratings in $\mathcal{D}_{pref}$. We use average entropy and variance as each instance in our dataset can be annotated by a different number of annotators, and there is no guarantee that the same annotator will be rating each of the mnemonics. We find that the average variance and entropy of our annotations are significantly lower than random chance.
Our datasets are based on publicly available GRE vocab words and mnemonics from MnemonicDictionary, and both were used as intended by the original authors. None of our datasets contain personal information, as users are referred to just by ID. All collected mnemonics are in English.
@@ B.4 Bayesian Model Evaluation
In this section, we evaluate the quality of our Hierarchical Bayesian model. Since our goal was to estimate the latent effectiveness of mnemonics, we first assess the convergence of our learned parameters across chains. All learned parameters have an $\hat{r}$ under 1.01 and an effective sample size over 1000, indicating strong convergence. Further, our final Bayesian preference label (which mnemonic has a higher effectiveness score) across chains has a Krippendorff’s $\alpha$ over 0.75, indicating strong agreement and convergence. Finally, in Figure 9, we display the log-likelihood values for our observed data across iterations, finding that they converge.
While convergence is more important to assess the quality of a Bayesian model that learns latent values, we also assess the generalizability of our model. We first train our model on 80% of our data and run inference on the remaining 20% for evaluation. In Figure 10, we compare the log-likelihood of predicting the observed data on the training and evaluation splits. For 3/5 of our observed data types, we find a non-significant difference between the log-likelihoods (2-sample $t$-test). The only significant difference is in the data associated with the learning preferences $y_{learn}$, further suggesting that modeling observed preferences is a challenging and interesting direction for future research.
@@ B.5 Obtaining Bayesian Preference Labels
To get the final latent overall mnemonic effectiveness values for training the DPO model $p_{dpo}(m|v)$, we average the latent variables over all five chains post-burn-in. For the ablation study where we compare optimizing on $y_{pair}$ versus $y_{bayes}$, we similarly obtain the final latent mnemonic effectiveness values by averaging the latent variables over all five chains post-burn-in, but this time we just use a random sample of the epochs (i.e. thinning). By taking a sample of each chain instead of using the entire chains, we introduce more variability and disagreement in the labels, allowing us to more meaningfully study the difference between optimizing on pairwise preferences versus all preference labels.
@@ C.1 Annotator Instructions
During our user studies, we ensure to provide ample instructions to annotators. On the home page of our flashcard learning app, users can view our Institutional Review Board documents, which detail the purpose of the user study and how user data will be collected and used. Further, we provide users instructions to help them rate the quality of mnemonic devices (Figure 7), which can be viewed at any time throughout the user study. All annotators are English speakers. Users were aware that they were participants in a research study and as participation was voluntary, compensation is fair.
@@ C.2 Qualitative Evaluation Details
In our qualitative evaluation, we compare our full model trained with DPO on Bayesian labels versus 10-shot GPT-4 (gpt-4-turbo-2024-04-09), where examples were chosen according to the highest latent quality scores in $\mathcal{D}_{train}$ (section 2.2). We also compare against expert human-written mnemonics to serve as an upper bound on mnemonic quality. These mnemonics were written by a professional copyeditor and creative writer with a Bachelor of Science degree that we hired on Upwork. As part of the interview, we asked the writer to produce two sample mnemonics to ensure the mnemonics would be high quality, and the annotator was paid a high rate of \$3 per mnemonic (around \$60/hr), which is fair for the participant demographic.
@@
C.3 Learning Time Distribution
We provide the distribution of learning time, measured in turns needed to recall the definition of the vocabulary term, for the winning and losing mnemonics in the pair. As expected, the winning mnemonics have a significantly lower ($p=0.05$) average number of turns needed until the definition is recalled, showing a clear gap in short-term learning efficacy between mnemonics in our pairs.
@@ D.2 Are Bayesian Tie-Breaks Good?
We previously found that using multiple preference labels to break ties within singular preference labels, improving LLM output quality with DPO (section 6.2). To ensure that these ties are better than random tie breaks, we use our GPT-4 classifier to compare the winning and losing mnemonics from our Bayesian labels (i.e. mnemonics with higher and lower effectiveness) where the majority vote in the pairwise setting is Tie. On the set of ties, GPT-4 states that our “winning” mnemonics is better 40% of the time, tied with the losing mnemonic 31% of the time, and is worse 29% of the time. Thus, even though humans marked these LLM outputs as tied, we were able to draw from other preference labels to identify a winning mnemonic, and GPT-4 also tends to agree that this mnemonic is higher quality.
@@ D.3 DPO Models versus Fine-Tuning
In Table 12, we use GPT-4 to judge the mnemonics produced by each of the DPO configurations used in section 6.2 versus the Smart model just using fine-tuning $p_{0}(m|v)$. We find that DPO improves the mnemonic quality of each of these models.
@@ D.4 Mnemonic Examples
In Table 13, we provide examples of high-quality keyword mnemonic devices generated by our final Smart model. We also show some examples of low-quality mnemonics from our model and highlight areas for improvement in Table 14.
@@
Figure 7: Instructions given to students to help them rate the quality of mnemonic devices for expressed preferences.
@@
Percentile Winning Mnemonic Turns Losing Mnemonic Turns
0 1.00 1.25
25 1.00 1.60
50 1.00 2.00
75 1.50 2.00
100 3.00 7.00
Average 1.24 2.13
Table 11: Distribution of learning time (Turns) for winning and losing mnemonics.